# DreamForge: Motion-Aware Autoregressive Video Generation for Multiview Driving Scenes

### v2

@@ Abstract
Recent advances in diffusion models have improved controllable streetscape generation and supported downstream perception and planning tasks. However, challenges remain in accurately modeling driving scenes and generating long videos. To alleviate these issues, we propose DreamForge, an advanced diffusion-based autoregressive video generation model tailored for 3D-controllable long-term generation. To enhance the lane and foreground generation, we introduce perspective guidance and integrate object-wise position encoding to incorporate local 3D correlation and improve foreground object modeling. We also propose motion-aware temporal attention to capture motion cues and appearance changes in videos. By leveraging motion frames and an autoregressive generation paradigm, we can autoregressively generate long videos (over 200 frames) using a 7-frame model, achieving superior quality compared to the baseline in 16-frame video evaluations. Finally, we integrate our method with the realistic simulation platform DriveArena to provide more reliable open-loop and closed-loop evaluations for vision-based driving agents. The project page is available at https://pjlab-adg.github.io/DriveArena/dreamforge.
@@ 1 Introduction
To enhance the diversity of driving scenes and facilitate downstream perception and planning tasks, recent approaches [15, 16, 17] have leveraged generative technologies such as NeRF [18], 3D GS [19], and diffusion models [20] to generate novel multiview driving scenes. Among these, diffusion-based methods [21, 17, 22, 23] have attracted significant attention for their ability to generate diverse, high-fidelity scenarios with flexible control conditions. However, these methods still encounter challenges, such as modeling geometrically and contextually accurate driving scenes and maintaining temporal coherence across long videos, which may affect their effectiveness in practical applications. On the other hand, these methods primarily use their pre-trained models for data augmentation in downstream tasks, and few methods [24, 13] involve the exploration of diffusion-based models for realistic generative simulations, which capture real-world visual and physical aspects, facilitate scalable scene generation, and support the ongoing development of AD algorithms within closed-loop systems.
To alleviate the above issues, following [17, 21], we design a diffusion-based framework, named DreamForge for multiview driving scene generation. Specifically, our DreamForge leverages flexible control conditions, e.g., road layouts and 3D bounding boxes, along with textual inputs, to generate geometrically and contextually accurate driving scenarios, maintaining cross-view and temporal consistency. By integrating perspective guidance, object-wise position encoding, and motion-aware autoregressive generation into conditional diffusion models [20, 25], our framework achieves significant improvements in several aspects: (1) Better controllability. We can not only control the generation of scenes with varying weather conditions and styles through texts, road layouts, and boxes but also improve street and foreground generation by perspective guidance (PG) and object-wise position encoding (OPE). The PG assists the network in learning to generate geometrically and contextually accurate driving scenes. The designed OPE enhances foreground modeling and naturally introduces local 3D correlation. (2) Better coherence. By learning motion cues from motion frame, ego pose, and feature differences and generating videos in an autoregressive manner, our DreamForge can generate multiview videos with flexible lengths using a model trained with short sequences while maintaining temporal coherence. Experiments demonstrate that we can generate long videos exceeding 200 frames using only a 7-frame model and achieve better generation quality than the baseline in 16-frame video evaluations.
@@ 2.2 Autonomous Driving Scene Generation
For driving scene generation, some studies [16, 35, 36, 15, 37] use NeRF [18] and 3D GS [19] for novel view synthesis by reconstructing scenes from logged videos, which often struggle with diverse weather and road layouts. On the other hand, recent advances in diffusion models [20] have established them as leading approaches [23, 22, 38, 21, 39, 40, 41, 42] for the generation of high-fidelity, diverse driving scenes through progressive denoising [43, 44]. For example, several methods [23, 45, 38] focused on the monocular diffusion-based world model, with ego actions to control ego-vehicle behavior and generate future scenes. DriveDreamer [34] and MagicDrive [17] employ HDmap and 3D box to enable more controllable scene generation. Recent methods, e.g., Panacea [21], DrivingDiffusion [22], and SubjectDrive [39], further advance 3D-controllable multiview video generation. Unlike these methods, we integrate perspective guidance, object-wise position encoding, and motion-aware autoregressive generation into diffusion models, resulting in significant improvements in both controllability and temporal coherence for long multiview video generation.
@@ 3 Methodology
3.1 Overview
We present our proposed DreamForge in Fig. 1 (a). Built upon the stable diffusion pipeline [25], DreamForge features diverse conditional encodings (e.g., text descriptions, camera poses, 3D bounding boxes and road layouts) to generate surround-view images with improved controllability (Sec. 3.2). To enhance street and foreground generation, we introduce perspective guidance and devise an object-wise position encoding (Sec. 3.2). To achieve long-term temporal coherence, we design a motion-aware temporal attention module and an autoregressive generation paradigm to facilitate seamless video generation (Sec. 3.3). Based on the enhanced controllability and temporal coherence, we integrate our DreamForge into a closed-loop simulation platform to demonstrate broader downstream applications (Sec. 3.4).
@@ 3.2 Improved Conditional Controllability
Conditional Encoding with ControlNet. Similar to MagicDrive [17], we utilize scene-level descriptions, camera poses, 3D bounding boxes of foreground objects, and the road layout of background elements as various forms of conditional encoding for controllable generation. Specifically, for scene-level encoding, we first enrich the text descriptions using GPT-4 and then utilize the CLIP text encoder [46] ($\mathcal{E}_{text}$) to extract the text embeddings $e_{text}$ from these descriptions. The camera poses $\{\mathbf{K}\in\mathbb{R}^{3\times 3},\mathbf{R}\in\mathbb{R}^{3\times 3},\mathbf{T}\in\mathbb{R}^{3\times 1}\}$ of each camera are encoded to $e_{cam}$ by Fourier Embedding [18] and MLP ($\mathcal{E}_{cam}$), where $\mathbf{K}$, $\mathbf{R}$, $\mathbf{T}$ represent camera intrinsic, rotations and translations respectively. For 3D boxes encoding, label embeddings are first extracted from the class labels using a text encoder. Coordinate embeddings are derived from the eight vertices of the 3D box through Fourier Embedding and MLP. Finally, both label and coordinate embeddings are combined and compressed into the final box embeddings $e_{box}$ using MLP. These embeddings, $e_{text}$, $e_{cam}$, and $e_{box}$, have the same dimensions and are concatenated before being fed into the ControlNet [47] and denoising U-Net, as shown in Fig. 1 (a). As for the road layout encoding, the 2D grid-formatted road layouts are processed through a ConvNet ($\mathcal{E}_{layout}$) to produce layout embeddings $e_{layout}$, which are then combined with the noised latents and fed into the ControlNet.
Perspective Guidance. As mentioned above, ControlNet encodes rich 3D information and camera poses, which could theoretically allow it to perform view transformation implicitly [17]; however, our experiments found that this implicit learning struggles to generate surround-view images that accurately align with the road layout, particularly in distant and complex areas, as illustrated in Fig. 2. Therefore, we further project the road layout and 3D boxes into the camera view using the camera poses to explicitly provide perspective guidance for position constraints and reduce the network’s difficulty in learning to generate geometrically and contextually accurate driving scenes. To this end, the contents of each category in the road layouts and the 3D boxes are projected onto the image plane of each camera to obtain the road canvas and box canvas, respectively. These canvases are concatenated to create the perspective canvas, which is encoded with the ConvNet ($\mathcal{E}_{canvas}$) to form the canvas embeddings $e_{canvas}$. The canvas embeddings are merged with the noised latents, and then input into ControlNet, as shown in Fig. 1 (a).
@@
Object-wise Position Encoding.
In addition to introducing a cross-view attention module [17, 21] that intuitively aggregates information globally from adjacent views to ensure multiview consistency, we also propose an object-wise position encoding to incorporate 3D object embeddings, enhancing foreground modeling and inherently providing local 3D correlation for improved object consistency, as shown in Fig. 5 and Appendix A.3. Following [48, 49], by using the transformation matrix of the camera of the $i$-th view, we transform the points $\textbf{P}_{i}^{c}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times 3}$ in the discrete camera frustum space of size $(W_{F},H_{F},D)$ into the common 3D world space for 3D coordinates $\textbf{P}^{3d}_{i}$. Subsequently, $\{\textbf{P}^{3d}_{i}\}_{i=1}^{N_{c}}$ from all $N_{c}$ camera views are aggregated and normalized to the range $[0,1]$ within the region of interest, producing normalized 3D points $\textbf{P}^{3d}\in\mathbb{R}^{N_{c}\times W_{F}\times H_{F}\times(D\times 3)}$. Then, we utilize the 3D boxes to generate the 3D masks $\textbf{M}^{3d}$, which indicate the foreground objects in points $\textbf{P}^{3d}$. Finally, the 3D points $\textbf{P}^{3d}$ and masks $\textbf{M}^{3d}$ are fed into the position encoder to obtain the 3D object position embeddings $\textbf{E}_{o}\in\mathbb{R}^{N_{c}\times W_{F}\times H_{F}\times C}$ of $N_{c}$ camera views. The position encoder is constructed with a stack of MLPs and aggregates the point features along the sampling ray. The detailed procedure can be expressed as:
@@
Augmented Spatial Attention. We further incorporate object-wise position encoding into the self-attention module of the denoising U-Net to enhance its capabilities of extracting object appearances. Specifically, given features $\textbf{Z}_{s}$, we add the embeddings $\textbf{E}_{o}$ to the $\textbf{Z}_{s}$ before feeding them into the self-attention for spatial attention. The operation of augmented spatial attention (ASA) is formulated as follows:
@@
Note that we reuse the self-attention layer of the denoising U-Net and only fine-tune the linear layer for query mapping of the self-attention module, thereby avoiding the introduction of additional parameters and resulting in minimal increases in computational overhead during inference.
@@ 3.3 Motion-aware Autoregressive Generation
Recent multiview driving scene generation works [17, 39] focus on fixed-length video generation but encounter challenges with extended videos due to memory limitations and poor temporal consistency. Some methods [21, 22] utilize keyframes as control conditions and enhance coherence through sliding windows; however, motion cues and temporal modeling remain inadequate, struggling to generate long videos. In this paper, we devise a motion-aware temporal attention module to improve consistency by incorporating motion cues from historical frames, ego poses, and feature differences, enabling our DreamForge to achieve effective autoregressive video generation.
@@
Let $\{\textbf{I}_{i}\}_{i=-M}^{-1}$ represent the $M$ motion frames sampled from the previous video clip. As shown in Fig. 1 (a), these motion frames are processed by the VAE encoder to extract motion latents, which are then fed into the U-Net using shared parameters with the denoising U-Net to generate multi-resolution motion features. During the denoising process, these motion features are concatenated with the corresponding noised latents to compute temporal attention. Additionally, we encode the relative poses between adjacent frames into the motion embedding for ego-motion cues and propose learning bidirectional local motion within the video clip to help the model understand changes in the background. Fig. 1 (c) illustrates the detailed architecture of the motion-aware temporal attention module (MTA). Specifically, given the motion features $\textbf{F}_{M}=\{\textbf{f}_{-M},...,\textbf{f}_{-1}\}\in\mathbb{R}^{HW\times M\times C}$ and the noised latents $\textbf{Z}_{T}=\{\textbf{z}_{0},...,\textbf{z}_{T-1}\}\in\mathbb{R}^{HW\times T\times C}$, where $M$, $T$, $H$, $W$, and $C$ denote the motion length, video length, spatial height, width, and number of channels, respectively, we calculate the motion-aware temporal attention as:
@@
To facilitate online inference and streaming video generation while maintaining temporal coherence, we employ an autoregressive generation pipeline. During inference, we randomly sample previously generated images as motion frames and calculate the corresponding relative ego poses to provide motion cues. This method enables the diffusion model to generate the current video clip with enhanced consistency, ensuring smoother transitions and improved coherence with previously generated frames. By utilizing motion frames, our method eliminates the need for a sliding window, thereby avoiding redundant generation. Our method also supports an optional post-processing strategy using sliding windows to further enhance temporal coherence between adjacent video clips. Please see the Appendix for more details.
@@ 4 Experiments
Our DreamForge is built on the pre-trained SD V1.5 [25]. The input resolution for each view is set to $224\times 400$ by default. The video length $T$ and length $M$ of motion frames are set to 7 and 2. Please see more details about implementation, datasets, and metrics in the Appendix A.1 and A.2; and find more video demos in the supplementary materials.
@@ 4.1 Quantitative Comparison
Table 2: Comparison of generation fidelity on generated 16-frame video clips from nuScene validation (tested by BEVFormer [55]). The frame “7/16” indicates training using 7 frames while inference with 16 frames.
@@
We present a video generation fidelity comparison in Table 2, evaluating the quality of the generated videos with a length of 16 frames. Compared to the baseline, i.e., temporal MagicDrive trained with 16-frame clips, our DreamForge achieves a lower FVD and significantly outperforms the baseline in terms of object mAP ($+3.51$ points) and map mIoU ($+10.73$ points). Furthermore, when upgrading the resolution to $336\times 600$, the object mAP shows a substantial increase, exceeding the baseline by 8.17 mAP. When the resolution continues to increase to $448\times 800$, the performance of the perception results improves further, but the FVD rises. This indicates that, while fixing the training iterations, a larger resolution is more beneficial for perception models, but it also complicates the similarity in synthetic data. Notably, we generate the required clips through motion-aware autoregressive generation using only our 7-frame model. Additionally, there is no need to retrain the model to produce longer videos, making our approach more efficient and resource-friendly.
@@
Table 6: Comparison of generation fidelity on generate images from nuScenes validation (tested by BEVFormer [55]).
@@
We further conduct experiments in DriveArena’s closed-loop mode. In this mode, the trajectory outputted by the driving agents at 2 Hz, consisting of six path points over the next 3 seconds, is interpolated to create a 10 Hz trajectory, which is then directly used for ego vehicle control. Without loss of generality, we perform closed-loop testing on two representative routes in Singapore-oneorth and Boston-seaport. PDM Score (PDMS) and Arena Drive Score (ADS) are evaluated, with detailed results presented in Table 5. From these results, we can also conclude that: (1) Upgrading DriveArena with our temporal version DreamForge boosts the PDMS scores for both the routes and the driving agents. (2) UniAD consistently outperforms VAD on these routes in the closed-loop mode of the upgraded DriveArena. (3) The performance of UniAD is more susceptible to temporal coherence, which aligns with the observations made in the open-loop mode. These findings demonstrate the temporal coherence brought by our DreamForge can facilitate the application of DriveArena for realistic autonomous driving simulations.

### v3

@@ Abstract
Recent advances in diffusion models have improved controllable streetscape generation and supported downstream perception and planning tasks. However, challenges remain in accurately modeling driving scenes and generating long videos. To alleviate these issues, we propose DreamForge, an advanced diffusion-based autoregressive video generation model tailored for 3D-controllable long-term generation. To enhance the lane and foreground generation, we introduce perspective guidance and design object-wise position encoding to incorporate local 3D correlation and improve foreground object modeling. We also propose motion-aware temporal attention to capture motion cues and appearance changes in videos. By leveraging motion frames and an autoregressive generation paradigm, we can autoregressively generate long videos (over 200 frames) using a model trained in short sequences, achieving superior quality compared to the baseline in 16-frame video evaluations. Finally, we integrate our method with the realistic simulator DriveArena to provide more reliable open-loop and closed-loop evaluations for vision-based driving agents. Project Page: https://pjlab-adg.github.io/DriveArena/dreamforge.
@@ 1 Introduction
To enhance the diversity of driving scenes and facilitate downstream perception and planning tasks, recent approaches [15, 16, 17] have leveraged generative technologies such as NeRF [18], 3D GS [19], and diffusion models [20] to generate novel multiview driving scenes. Among these, diffusion-based methods [21, 17, 22, 23, 24] have attracted significant attention for their ability to generate diverse, high-fidelity scenarios with flexible control conditions. However, these methods still encounter challenges, such as modeling geometrically and contextually accurate driving scenes and maintaining temporal coherence across long videos, which may affect their effectiveness in practical applications. On the other hand, these methods primarily use their pre-trained models for data augmentation in downstream tasks, and few methods [25, 13, 26] involve the exploration of diffusion-based models for realistic generative simulations, which capture real-world visual and physical aspects, facilitate scalable scene generation, and support the ongoing development of AD algorithms within closed-loop systems.
To alleviate the above issues, following [17, 21], we design a diffusion-based framework, named DreamForge for multiview driving scene generation. Specifically, our DreamForge leverages flexible control conditions, e.g., road layouts and 3D bounding boxes, along with textual inputs, to generate geometrically and contextually accurate driving scenarios, maintaining cross-view and temporal consistency. By integrating perspective guidance, object-wise position encoding, and motion-aware autoregressive generation into conditional diffusion models [20, 27], our framework achieves significant improvements in several aspects: (1) Better controllability. We can not only control the generation of scenes with varying weather conditions and styles through texts, road layouts, and boxes but also improve street and foreground generation by perspective guidance (PG) and object-wise position encoding (OPE). The PG assists the network in learning to generate geometrically and contextually accurate driving scenes. The designed OPE enhances foreground modeling and naturally introduces local 3D correlation. (2) Better coherence. By learning motion cues from motion frame, ego pose, and feature differences and generating videos in an autoregressive manner, our DreamForge can generate multiview videos with flexible lengths using a model trained with short sequences while maintaining temporal coherence. Experiments demonstrate that we can generate long videos exceeding 200 frames using only a model trained in short sequences and achieve better generation quality than the baseline in 16-frame video evaluations. In particular, our proposed DreamForge can adapt to various generative base models, such as SD V1.5 [27] and DiT [28], demonstrating its broader application to the autonomous driving community.
@@ 2.2 Autonomous Driving Scene Generation
For driving scene generation, some studies [16, 38, 39, 15, 40] use NeRF [18] and 3D GS [19] for novel view synthesis by reconstructing scenes from logged videos, which often struggle with diverse weather and road layouts. On the other hand, recent advances in diffusion models [20] have established them as leading approaches [23, 22, 41, 21, 42, 43, 44, 45, 46, 47] for the generation of high-fidelity, diverse driving scenes through progressive denoising [48, 49]. For example, several methods [23, 50, 41] focused on the monocular diffusion-based world model, with ego actions to control ego-vehicle behavior and generate future scenes. DriveDreamer [37] and MagicDrive [17] employ HDmap and 3D box to enable more controllable scene generation. Recent methods, e.g., Panacea [21], DrivingDiffusion [22], and SubjectDrive [42], further advance 3D-controllable multiview video generation. Unlike these methods, we integrate perspective guidance, object-wise position encoding, and motion-aware autoregressive generation into diffusion models, resulting in significant improvements in both controllability and temporal coherence for long multiview video generation.
@@ 3 Methodology
We present our proposed DreamForge in Fig. 1 (a). It features diverse conditional encodings for improved controllability, perspective guidance, and object-wise position encoding for enhanced street and foreground generation (Sec.3.1). Also, a motion-aware temporal attention module and autoregressive generation are designed to enable seamless video generation (Sec.3.2), allowing integration into a simulation platform for broader applications (Sec.3.3).
@@ 3.1 Improved Conditional Controllability
Conditional Encoding with ControlNet. Similar to MagicDrive [17], we utilize scene-level descriptions, camera poses, 3D bounding boxes of foreground objects, and the road layout of background elements as various forms of conditional encoding for controllable generation. Specifically, for scene-level encoding, we first enrich the text descriptions using GPT-4 and then utilize the CLIP text encoder [51] ($\mathcal{E}_{text}$) to extract the text embeddings $e_{text}$ from these descriptions. The camera poses $\{\mathbf{K}\in\mathbb{R}^{3\times 3},\mathbf{R}\in\mathbb{R}^{3\times 3},\mathbf{T}\in\mathbb{R}^{3\times 1}\}$ of each camera are encoded to $e_{cam}$ by Fourier Embedding [18] and MLP ($\mathcal{E}_{cam}$), where $\mathbf{K}$, $\mathbf{R}$, $\mathbf{T}$ represent camera intrinsic, rotations and translations respectively. For 3D boxes encoding, label embeddings are first extracted from the class labels using a text encoder. Coordinate embeddings are derived from the eight vertices of the 3D box through Fourier Embedding and MLP. Finally, both label and coordinate embeddings are combined and compressed into the final box embeddings $e_{box}$ using MLP. These embeddings, $e_{text}$, $e_{cam}$, and $e_{box}$, have the same dimensions and are concatenated before being fed into the ControlNet [52] and denoising blocks, as shown in Fig. 1 (a). As for the road layout encoding, the 2D grid-formatted road layouts are processed through a ConvNet ($\mathcal{E}_{layout}$) to produce layout embeddings $e_{layout}$, which are then combined with the noised latents and fed into the ControlNet.
Perspective Guidance. As mentioned above, ControlNet encodes rich 3D information and camera poses, which could theoretically allow it to perform view transformation implicitly [17]; however, our experiments found that this implicit learning struggles to generate surround-view images that accurately align with the road layout, particularly in distant and complex areas, as illustrated in Fig. 2. Therefore, we further project the road layout and 3D boxes into the camera view using the camera poses to explicitly provide perspective guidance for position constraints and reduce the network’s difficulty in learning to generate geometrically and contextually accurate driving scenes. To this end, the contents of each category in the road layouts and the 3D boxes are projected onto the image plane of each camera to obtain the road canvas and the box canvas, respectively. Specifically, we plot the contents of each category on a dedicated channel, encoding road and box classes in one-hot maps where 1 indicates the presence of the category and 0 otherwise. This approach enforces clear and structured constraints on the perspective view, enabling the model to effectively generate diverse elements in complex scenes. Subsequently, these canvases are concatenated to create the perspective canvas, which is encoded with ConvNet ($\mathcal{E}_{canvas}$) to form the canvas embeddings $e_{canvas}$. The canvas embeddings are merged with the noised latents, and then input into ControlNet, as shown in Fig. 1 (a).
@@
Object-wise Position Encoding. In addition to introducing a cross-view attention module [17, 21] that intuitively aggregates information globally from adjacent views to ensure multiview consistency, we also propose an object-wise position encoding to incorporate 3D object embeddings, enhancing foreground modeling and inherently providing local 3D correlation for improved object consistency, as shown in Fig. 5 and Appendix A.4.2. Following [53, 54], by using the transformation matrix of the camera of the $i$-th view, we transform the points $\textbf{P}_{i}^{c}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times 3}$ in the discrete camera frustum space of size $(W_{F},H_{F},D)$ into the common 3D world space for 3D coordinates $\textbf{P}^{3d}_{i}$. Subsequently, $\{\textbf{P}^{3d}_{i}\}_{i=1}^{N_{c}}$ from all $N_{c}$ camera views are aggregated and normalized to the range $[0,1]$ within the region of interest, producing normalized 3D points $\textbf{P}^{3d}\in\mathbb{R}^{N_{c}\times W_{F}\times H_{F}\times(D\times 3)}$. Then, we utilize the 3D boxes to generate the 3D masks $\textbf{M}^{3d}$, which indicate the foreground objects in points $\textbf{P}^{3d}$. Finally, the 3D points $\textbf{P}^{3d}$ and masks $\textbf{M}^{3d}$ are fed into the position encoder to obtain the 3D object position embeddings $\textbf{E}_{o}\in\mathbb{R}^{N_{c}\times W_{F}\times H_{F}\times C}$ of $N_{c}$ camera views. The position encoder is constructed with a stack of MLPs and aggregates the point features along the sampling ray. The detailed procedure can be expressed as:
@@
Augmented Spatial Attention. We further incorporate object-wise position encoding into the self-attention module of the denoising blocks to enhance its capabilities of extracting object appearances. Specifically, given features $\textbf{Z}_{s}$, we add the embeddings $\textbf{E}_{o}$ to the $\textbf{Z}_{s}$ before feeding them into the self-attention for spatial attention. The operation of augmented spatial attention (ASA) is formulated as follows:
@@
Note that we reuse the self-attention layer of the denoising blocks and only fine-tune the linear layer for query mapping of the self-attention module, thereby avoiding the introduction of additional parameters and resulting in minimal increases in computational overhead during inference.
@@ 3.2 Motion-aware Autoregressive Generation
Recent multiview driving scene generation works [17, 42] focus on fixed-length video generation but face challenges with extended videos due to memory limitations and poor temporal consistency. Some methods [21, 22] use keyframes as control conditions and enhance coherence through sliding windows, yet motion cues and temporal modeling remain insufficient for long video generation. We design the motion-aware temporal attention to improve consistency by integrating motion cues from historical frames, ego poses, and feature differences, enabling our method to achieve effective autoregressive video generation.
@@
Let $\{\textbf{I}_{i}\}_{i=-M}^{-1}$ represent the $M$ motion frames sampled from the previous video clip. As shown in Fig. 1 (a), these motion frames are processed by the VAE encoder to extract motion latents, which are then fed into the blocks using shared parameters with the denoising blocks to generate multi-resolution motion features. During the denoising process, these motion features are concatenated with the corresponding noised latents to compute temporal attention. Additionally, we encode the relative poses between adjacent frames into the motion embedding for ego-motion cues and propose learning bidirectional local motion within the video clip to help the model understand changes in the background. Fig. 1 (c) illustrates the detailed architecture of the motion-aware temporal attention module (MTA). Specifically, given the motion features $\textbf{F}_{M}=\{\textbf{f}_{-M},...,\textbf{f}_{-1}\}\in\mathbb{R}^{HW\times M\times C}$ and the noised latents $\textbf{Z}_{T}=\{\textbf{z}_{0},...,\textbf{z}_{T-1}\}\in\mathbb{R}^{HW\times T\times C}$, where $M$, $T$, $H$, $W$, and $C$ denote the motion length, video length, spatial height, width, and number of channels, respectively, we calculate the MTA as:
@@
To facilitate online inference and streaming video generation while maintaining temporal coherence, we employ an autoregressive generation pipeline. During inference, we sample previously generated images as motion frames and calculate the corresponding relative ego poses to provide motion cues. This method enables the diffusion model to generate the current video clip with enhanced consistency, ensuring smoother transitions and improved coherence with previously generated frames. By utilizing motion frames, our method eliminates the need for a sliding window, thereby avoiding redundant generation. Our method also supports an optional post-processing strategy using sliding windows to further enhance temporal coherence between adjacent video clips. Please see the Appendix A.2.1 for more details.
@@ 4 Experiments
Our DreamForge is built on SD V1.5 [56] by default. We also offer a version using DiT [28] and 3D VAE [57] as the base model, as detailed in Sec.4.3 and Table 3. The default input resolution for each view is $224\times 400$, with a video length $T$ of 7 and motion frames length $M$ of 2. Please see more details about implementation, datasets, and metrics in the Appendix; and find more demos on project page.
@@ 4.1 Quantitative Comparison
Table 2: Comparison of generation fidelity on generated 16-frame clips from nuScene validation (tested by BEVFormer [62]). “7/16” indicates training using 7 frames while inference with 16 frames.
Data Source Resolution Base Model FVD $\downarrow$ mAP $\uparrow$ mIoU $\uparrow$
Ori nuScenes 224$\times$400 - - 29.69 36.70
MagicDrive [17] 224$\times$400 SD V1.5 218.12 11.86 18.34
MagicDrive3D [63] 224$\times$400 SD V1.5 210.40 12.05 18.24
DreamForge 224$\times$400 SD V1.5 209.90 14.37 29.07
DreamForge 448$\times$800 SD V1.5 233.20 22.52 32.98
MagicDriveDiT [64] 848$\times$1600 3DVAE, DiT 94.84 18.17 20.40
DreamForge 448$\times$800 3DVAE, DiT 103.61 19.17 34.36
Table 3: Comparison of the model under different base model configurations. Metrics are computed for 16-frame video clips.
@@
We present a video generation fidelity comparison in Table 2, evaluating the quality of the generated videos with a length of 16 frames. Compared to the baseline, i.e., temporal MagicDrive trained with 16-frame clips, our DreamForge achieves a lower FVD and significantly outperforms the baseline in terms of object mAP ($+3.51$ points) and map mIoU ($+10.73$ points). Furthermore, when upgrading the resolution to $336\times 600$, the object mAP shows a substantial increase, exceeding the baseline by 8.17 mAP. When the resolution continues to increase to $448\times 800$, the performance of the perception results improves further, but the FVD rises. This indicates that, while fixing the training iterations, a larger resolution is more beneficial for perception models, but it also complicates the similarity in synthetic data. Notably, we generate the required clips through motion-aware autoregressive generation using only our model trained in short sequences (e.g. 7 frames). There is no need to retrain the model to produce longer videos, making our approach more efficient and resource-friendly.
@@
Table 7: Comparison of generation fidelity on generate images from nuScenes validation (tested by BEVFormer [62]). “PG”, “OPE” and “ASA” denote perspective guidance, object-wise position encoding, and augmented spatial attention, respectively.
@@
We further conduct experiments in DriveArena’s closed-loop mode. In this mode, the trajectory outputted by the driving agents at 2 Hz, consisting of six path points over the next 3 seconds, is interpolated to create a 10 Hz trajectory, which is then directly used for ego vehicle control. Without loss of generality, we perform closed-loop testing on two representative routes in Singapore-oneorth and Boston-seaport. PDM Score (PDMS) and Arena Drive Score (ADS) are evaluated, with detailed results presented in Table 6. From these results, we can also conclude that: (1) Upgrading DriveArena with our temporal version DreamForge boosts the PDMS scores for both the routes and the driving agents. (2) UniAD consistently outperforms VAD on these routes in the closed-loop mode of the upgraded DriveArena. (3) The performance of UniAD is more susceptible to temporal coherence, which aligns with the observations made in the open-loop mode. These findings demonstrate the temporal coherence brought by our DreamForge can facilitate the application of DriveArena for realistic simulations.